a net of a solid shape . The two circular faces are congruent and each has a diameter of 7cm nd a hight of 25cm. Calculate the total surface are nd the volume of the solid shape formed
Answer by addingup(3677)

net of a solid shape with two PARALLEL congruent circular faces. The shape has a height of 25 and each circular face has a diameter of 7.
:
Diameter 7, radius = D/2 = 7/2 = 3.5
Height: 25
-------------------------------
Area of the circle:
Pir^2 = A
Pi*3.5^2 = 38.48 this is the area of each of the circular faces, and we have 2 faces: 38.48*2 = 76.96
-------------------
:Area of the cylinder:
Perimeter of the faces:
dPi = P
7*Pi = 22
Area of the side:
22*25 = 550 Now add the area of the faces:
550+76.95 = 626.96 you may have to round off to 627